# Towards Fast and Scalable Normal Integration using Continuous Components

### v1

@@ 1 Introduction
The problem of reconstructing a surface from its normal map, known as normal integration, arises naturally in many applications of computer vision, particularly in photometric shape recovery, such as shape from shading [12] and from polarization [1, 2] and photometric stereo [25, 13], in which the objective is to reconstruct a 3D surface from shading information, using estimated normals as an intermediate step.
@@
The mathematical problem underlying normal estimation has been extensively studied and several recent methods have been proposed that achieve sub-millimeter accuracy on object-level normal maps [5, 16, 19]. While accurate, state-of-the-art approaches are predominantly based on a costly global optimization that jointly estimates the depth value at each pixel, with the number of optimization variables and constraints thereby scaling with the number of pixels. Additionally, multiple optimization steps with iterative reweighting are necessary to capture surface discontinuities, which cannot be correctly estimated a priori in the general case and would otherwise cause the reconstruction to be suboptimal [5]. As a consequence, these methods have runtimes in the order of minutes for relatively low scales (typically in the order of $10^{4}$ valid pixels) and up to hours for high resolutions and scales.
@@
To address the above limitation, we observe that naturally occurring surfaces are often made of several smooth regions. If each such region was independently reconstructed, obtaining a single, global surface would reduce to optimally aligning each region to one another. Thus, we propose to recast normal integration into the estimation of the relative scales of continuous surface components. We introduce a simple and effective heuristic, based on normal similarity, to identify continuous components, and independently reconstruct each of them using the state-of-the-art formulation of [19]. We then jointly scale the depth of all pixels in each component by optimizing a single scale parameter for each component. For this, we reframe [19] and the discontinuity model of BiNI [5] to use relative component scales as optimization variables, and additionally introduce an outlier reweighting mechanism that rebalances the optimization terms. Importantly, we find that this reweighting significantly speeds up convergence also for previous, pixel-level methods, although their scalability remains limited due to the size of their optimization problem. Through its component-based formulation, our method greatly reduces the number of variables and constraints used in the optimization, resulting in a reduction of one order of magnitude in the execution time for mid-to-high resolution normal maps. Additionally, our approach achieves state-of-the-art reconstruction accuracy on the DiLiGenT benchmark [22] for normal integration in as little as a few seconds.
@@
Thus, our contribution is a framework that recasts normal integration as an estimation of relative scales of continuous components, which (i) achieves state-of-the-art reconstruction accuracy on normal integration benchmarks, and (ii) enables scaling discontinuity-preserving normal integration to larger resolutions with an order of magnitude reduction in the execution time. Our code will be made open source.
@@ 2 Related work
We refer the reader to [20, 21] for extensive surveys of methods for normal integration. In the following Section, we focus on the most recent, state-of-the-art approaches.
A key challenge in normal integration is that the input normal map might represent a surface with discontinuities, which naturally arise, for instance due to self-occlusions. Failure to preserve such discontinuities results in overly continuous reconstructions with global distortions [28, 4]. To address this problem, a number of single-analysis methods have proposed detecting discontinuities as a preprocessing step, according to different strategies [15, 26, 24, 27]. These methods, however, tend not to be robust, since errors in the initial discontinuity detection cannot be later corrected during the reconstruction process. Recently, large improvements have been achieved by methods that estimate discontinuities through an iteratively reweighted optimization problem [5, 15, 19]. While accurate, these methods suffer from long execution times and scale poorly to larger resolutions, due to the need to jointly recover a depth value at all pixels through a single, global optimization problem.
To address this problem, Heep and Zell [10] have recently proposed a meshing-based approach for normal integration, in which a triangle mesh is precomputed from the input normal map based on estimated curvature and a surface reconstruction is obtained by optimizing the depth of each mesh vertex. While greatly reducing the number of variables and the execution time, the single, fixed mesh representation does not support modeling discontinuities, which is admittedly not straightforward and would require operations to realign the mesh and adjust its topology [10]. In this work, we aim at achieving the best of both discontinuity-preserving and computationally efficient approaches. Our component-based formulation is naturally compatible with state-of-the-art pixel-level methods, but also allows effectively preserving discontinuities while significantly improving scalability.
@@ 3.1 Background
Surface normal integration is the problem of reconstructing a depth map from an input surface normal map, assuming known camera intrinsic parameters. Using the notation of [19], for a generic pair of neighboring pixels $(a,b)$ with pixel coordinates $(u_{a},v_{a})^{\mathsf{T}}$ and $(u_{b},v_{b})^{\mathsf{T}}$, respectively, we denote with $\boldsymbol{n_{a}}=\left(n_{ax},n_{ay},n_{az}\right)^{\mathsf{T}}$ and $\boldsymbol{n_{b}}=\left(n_{bx},n_{by},n_{bz}\right)^{\mathsf{T}}$ the surface normal vectors at those pixels. The values of the depth $z_{a}$, $z_{b}$ at these pixels can then be modeled through a linear relationship in logarithmic space, which we refer to as a model of continuity:
@@
where $\tilde{z}_{a}\coloneqq\log(z_{a})$, $\tilde{z}_{b}\coloneqq\log(z_{b})$, and the coefficient $\tilde{\omega}_{b\rightarrow a}$ assumes a different value depending on the formulation. For instance, in BiNI [5], $\tilde{\omega}_{b\rightarrow a}$ has the following form for a pinhole camera of focal lengths $f_{x}$ and $f_{y}$ and principal point $(c_{x},c_{y})$:
@@
Milano et al[19] instead derive the following alternative equation for generic central cameras:
@@
Figure 1: Example of surface discontinuities. Left: The chair in the foreground is separated by a full surface discontinuity from the wall in the background. Infinitely many depth maps can describe the input normal map, and solutions of the integration are up to the relative scales of the chair and the wall. Right: The surface points $\boldsymbol{p_{a}}$ and $\boldsymbol{p_{b}}$ are separated by a discontinuity along the path in red, but the yellow path connects them along the surface without discontinuities. Source of object meshes : [17] (left), [22] (right).
While in the general case the integration solution is up to these residuals, a number of assumptions can be made about the discontinuities. In particular, discontinuities often arise from self-occlusions rather than from fully disconnected surfaces, for instance, in normal maps that capture a single object [22]. In general, even in the presence of multiple objects, while the residual $\chi_{b\rightarrow a}$ between two neighboring pixels $a$ and $b$ might be significant in magnitude, there often exists an alternative path along the surface that connects the two corresponding surface points along a trajectory with no discontinuities (Fig.˜1, right).
State-of-the-art methods implicitly exploit this fact by assigning a lower weight in the optimization objective to the constraints (1) that, over the course of the optimization, are found to have a large residual magnitude. The rationale is that if all constraints were equally weighted, those associated with discontinuities would incorrectly drive the optimization to assign them a small residual, causing the reconstruction to converge to an overly continuous surface. By iteratively adapting the weights of the constraints, these methods instead focus on the more continuous paths along the surface, and thereby more accurately recover discontinuities, while still reconstructing a connected surface. A complementary approach, recently proposed by [19], consists in dynamically updating the coefficients $\tilde{\omega}_{b\rightarrow a}$ in the equations (1) that have large residuals, so as to explicitly take into account the magnitude of the discontinuities. This is achieved through an additive term in the logarithm (3), which admits a geometrical interpretation. We refer to the exact mechanism by which the above methods achieve the reweighting of the constraints as a model of discontinuity.
The leading model of discontinuity is the bilateral weighting scheme proposed by BiNI [5] which acts through two different mechanisms. First, it relies on the assumption that surfaces are semi-smooth. This implies that, indexing the two neighboring pixels $b$ and b-b on opposite sides of a pixel $a$, the depth map has a discontinuity between at most one of the sides $(a,b)$ or (a,b)(a,-b). This is modeled by weighting the constraint (1) between $a$ and $b$ by the factor,
@@
As noted by [19], the factor $\gamma_{b\rightarrow a}$ can be expressed as
@@
where $f$ is the focal length of the camera, and is crucial to the optimization convergence, which it affects through (6) by introducing a multiplicative factor $(\boldsymbol{n_{a}}^{\mathsf{T}}\boldsymbol{\tau_{a}})^{2}$. Since the quantity $\boldsymbol{n_{a}}^{\mathsf{T}}\boldsymbol{\tau_{a}}$ tends to 0 for pixels $a$ that lie close to an object boundary [18], this weighting effectively assigns lower confidence to constraints involving pixels that are likely to lie next to a depth discontinuity.
@@ 3.3 Formation and filling of the initial components
Importantly, since no two components share any variables or constraints, the above optimization problem can be solved independently and in parallel for each of the components. Similarly to previous log-depth-based methods [5, 19], we initialize all log-depth values to 0 and perform optimization using the conjugate-gradient method [11], denoted as $\mathrm{cg}(\mathbf{A},\mathbf{b})$, on the normal equations:
@@ 4.1 Experimental settings
Our implementation relies on a combination of GPU-accelerated tensor operations and CPU-based operations (e.g., conjugate-gradient-based optimization, for which we use the SciPy library [23]). We compare our method to the current state-of-the-art approach of [19], noting that we also use their model of continuity in our framework. We reimplement the baseline and integrate it into our framework, to decouple the effects of GPU acceleration and formulation-specific computational aspects. All experiments are run on a single NVIDIA RTX $3080$ Laptop GPU.
@@ 4.2 Benchmark experiments
We evaluate on the DiLiGenT benchmark [22] for normal integration, which features ground-truth, object-level normal maps of resolution $608\times 512$. In these maps the background is masked out, and a normal vector is defined only at the object pixels, resulting in 24 70624\,706 to 56 56056\,560 depth values to be estimated. For each object and run we report the mean average depth error (MADE) and the total execution time.
@@
For both methods we use $8$-connectivity and convergence parameters $\Delta E_{\mathrm{max}}=10^{-3}$, $T=150$. We refer the reader to the Supplementary for ablations on these parameters. Additionally, while slight speed-ups of our method can be achieved through our optional merging (see Supplementary), we find that this effect is not particularly significant, given the small number of valid pixels in the input normal maps. We therefore do not use merging in our DiLiGenT experiments, and show instead its benefit on larger-resolution maps in the next Section. We test our method for different values of the normal similarity threshold $\theta_{c}$, including the limit case in which each pixel is assigned to a component, which we denote as $\theta_{c}=\mathrm{None}$. For the baseline, we include both the version with and without discontinuity computation ($\alpha_{b\rightarrow a}$ in [19]). Finally, we note that our proposed outlier reweighting mechanism can also be applied to the log-depth-based optimization of [19] and we therefore evaluate both methods with and without it.
@@
The results of our evaluation are shown in Table˜1. We start by noting that our outlier reweighting significantly improves the convergence time for pixel-level approaches, not only for our relative scale optimization, but also for the log-depth-based baseline of [19]. Importantly, for this baseline, outlier reweighting also improved accuracy, although with the exception of some objects (cf., buddha, cow, pot1). Additionally, we note that while the discontinuity computation strategy of [19] is beneficial for objects with large discontinuities (buddha, harvest, $\mathrm{D}$ vs. $\mathrm{ND}$ in Table˜1), its effect is partially reduced by the use of outlier reweighting.
@@
Crucially, for our component-level versions, outlier reweighting is critical for achieving optimal accuracy, resulting in significantly better and more stable convergence across the board, particularly for objects with large discontinuities (buddha, harvest, reading, goblet), as can be seen by comparing successive rows in Table˜1. Overall, while with outlier reweighting the performance of the two methods is comparable for the most continuous objects, our component-based variants achieve better accuracy than log-depth optimization for the more discontinuous objects (buddha, harvest, reading). This effect is a function of the chosen value for the threshold $\theta_{c}$, with smaller values generally resulting in lower error at the cost of slightly increased runtime. We refer to the Supplementary for details on the minimum theoretical error that can be achieved for different values of $\theta_{c}$. Notably, our pixel-level version also slightly outperforms the (pixel-level) formulation of [19], suggesting that solving for relative scales instead of log-depth may lead to a more well-posed optimization problem.
@@
Finally, we note that, when using outlier reweighting both our component-based formulation and the pixel-level method of [19] achieve comparable runtime on the DiLiGenT benchmark. As we show in the next Section, however, the gap between the execution times of the two frameworks becomes significantly larger as the number of valid pixels in the normal map increases.
@@
[19], $\mathrm{D}$ N/A $\mathbf{0.02}$ $2.70$ $0.62$ $3.32$ $0.05$ $2.98$ $0.13$ $1.49$ $2.07$ $6.96$ $0.47$ $\underline{3.74}$ $0.15$ $2.86$ $0.25$ $2.02$ $\mathbf{4.04}$ $3.45$
@@
[19], $\mathrm{ND}$ N/A $\mathbf{0.02}$ $1.84$ $0.70$ $3.04$ $0.05$ $1.87$ $0.13$ $1.15$ $2.21$ $5.99$ $0.47$ $4.13$ $\underline{0.14}$ $\mathbf{1.79}$ $0.24$ $\underline{1.62}$ $\mathbf{4.04}$ $1.62$
@@
Table 1: Mean absolute depth error (MADE, abbreviated as $\mathrm{Err}$) $\mathrm{m}\mathrm{m}$] and total execution time (abbreviated as $t$) $\mathrm{s}$] on the DiLiGenT benchmark [22]. For each object, bold and underlined denote respectively the best and the second-best result across the methods. All experiments use $\Delta E_{\mathrm{max}}=10^{-3}$, $T=150$, and 88-connectivity, without merging. $\mathrm{D}$ and $\mathrm{ND}$ denote the version of [19] with and without discontinuity computation, respectively ($\alpha_{b\rightarrow a}$ in [19]). *This object contains a full depth discontinuity.

### v2

@@ 1 Introduction
The problem of reconstructing a surface from its normal map, known as normal integration, arises naturally in many applications of computer vision, particularly in photometric shape recovery, such as shape from shading [Horn1975ShapeFromShading] and from polarization [Atkinson2006RecoverySurfaceOrientation, Ba2020DeepShapeFromPolarization] and photometric stereo [Woodham1980PhotometricStereo, Ikehata2023SDMUniPS], in which the objective is to reconstruct a 3D surface from shading information, using estimated normals as an intermediate step.
@@
The mathematical problem underlying normal estimation has been extensively studied and several recent methods have been proposed that achieve sub-millimeter accuracy on object-level normal maps [Cao2022BiNI, Kim2024DiscontinuityPreserving, Milano2025DiscontinuityAwareNormalIntegration]. While accurate, state-of-the-art approaches are predominantly based on a costly global optimization that jointly estimates the depth value at each pixel, with the number of optimization variables and constraints thereby scaling with the number of pixels. Additionally, multiple optimization steps with iterative reweighting are necessary to capture surface discontinuities, which cannot be correctly estimated a priori in the general case and would otherwise cause the reconstruction to be suboptimal [Cao2022BiNI]. As a consequence, these methods have runtimes in the order of minutes for relatively low scales (typically in the order of $10^{4}$ valid pixels) and up to hours for high resolutions and scales.
@@
To address the above limitation, we observe that naturally occurring surfaces are often made of several smooth regions. If each such region was independently reconstructed, obtaining a single, global surface would reduce to optimally aligning each region to one another. Thus, we propose to recast normal integration into the estimation of the relative scales of continuous surface components. We introduce a simple and effective heuristic, based on normal similarity, to identify continuous components, and independently reconstruct each of them using the state-of-the-art formulation of [Milano2025DiscontinuityAwareNormalIntegration]. We then jointly scale the depth of all pixels in each component by optimizing a single scale parameter for each component. For this, we reframe [Milano2025DiscontinuityAwareNormalIntegration] and the discontinuity model of BiNI [Cao2022BiNI] to use relative component scales as optimization variables, and additionally introduce an outlier reweighting mechanism that rebalances the optimization terms. Importantly, we find that this reweighting significantly speeds up convergence also for previous, pixel-level methods, although their scalability remains limited due to the size of their optimization problem. Through its component-based formulation, our method greatly reduces the number of variables and constraints used in the optimization, resulting in a reduction of one order of magnitude in the execution time for mid-to-high resolution normal maps. Additionally, our approach achieves state-of-the-art reconstruction accuracy on the DiLiGenT benchmark [Shi2016DiLiGenT] for normal integration in as little as a few seconds.
@@
Thus, our contribution is a framework that recasts normal integration as an estimation of relative scales of continuous components, which (i) achieves state-of-the-art reconstruction accuracy on normal integration benchmarks, and (ii) enables scaling discontinuity-preserving normal integration to larger resolutions with an order of magnitude reduction in the execution time. Our code is publicly available [^1] https://github.com/francescomilano172/normal_integration_continuous_components..
@@ 2 Related work
We refer the reader to [Queau2018NormalIntegrationSurvey, Queau2018VariationMethodsNormalIntegration] for extensive surveys of methods for normal integration. In the following Section, we focus on the most recent, state-of-the-art approaches.
A key challenge in normal integration is that the input normal map might represent a surface with discontinuities, which naturally arise, for instance due to self-occlusions. Failure to preserve such discontinuities results in overly continuous reconstructions with global distortions [Zhu2020LeastSquaresSurfaceReconstruction, Cao2021NormalIntegrationInversePlaneFitting]. To address this problem, a number of single-analysis methods have proposed detecting discontinuities as a preprocessing step, according to different strategies [Karacali2003ReconstructingDiscontinuous, Wu2006VisibleSurfaceReconstruction, Wang2012DetectingDiscontinuitiesSurfaceReconstruction, Xie2019ARobustDGPBased]. These methods, however, tend not to be robust, since errors in the initial discontinuity detection cannot be later corrected during the reconstruction process. Recently, large improvements have been achieved by methods that estimate discontinuities through an iteratively reweighted optimization problem [Cao2022BiNI, Karacali2003ReconstructingDiscontinuous, Milano2025DiscontinuityAwareNormalIntegration]. While accurate, these methods suffer from long execution times and scale poorly to larger resolutions, due to the need to jointly recover a depth value at all pixels through a single, global optimization problem.
To address this problem, Heep and Zell [Heep2024AdaptiveScreenSpaceMeshingApproach] have recently proposed a meshing-based approach for normal integration, in which a triangle mesh is precomputed from the input normal map based on estimated curvature and a surface reconstruction is obtained by optimizing the depth of each mesh vertex. While greatly reducing the number of variables and the execution time, the single, fixed mesh representation does not support modeling discontinuities, which is admittedly not straightforward and would require operations to realign the mesh and adjust its topology [Heep2024AdaptiveScreenSpaceMeshingApproach]. In this work, we aim at achieving the best of both discontinuity-preserving and computationally efficient approaches. Our component-based formulation is naturally compatible with state-of-the-art pixel-level methods, but also allows effectively preserving discontinuities while significantly improving scalability.
@@ 3.1 Background
Surface normal integration is the problem of reconstructing a depth map from an input surface normal map, assuming known camera intrinsic parameters. Using the notation of [Milano2025DiscontinuityAwareNormalIntegration], for a generic pair of neighboring pixels $(a,b)$ with pixel coordinates $(u_{a},v_{a})^{\mathsf{T}}$ and $(u_{b},v_{b})^{\mathsf{T}}$, respectively, we denote with $\boldsymbol{n_{a}}=\left(n_{ax},n_{ay},n_{az}\right)^{\mathsf{T}}$ and $\boldsymbol{n_{b}}=\left(n_{bx},n_{by},n_{bz}\right)^{\mathsf{T}}$ the surface normal vectors at those pixels. The values of the depth $z_{a}$, $z_{b}$ at these pixels can then be modeled through a linear relationship in logarithmic space, which we refer to as a model of continuity:
@@
where $\tilde{z}_{a}\coloneqq\log(z_{a})$, $\tilde{z}_{b}\coloneqq\log(z_{b})$, and the coefficient $\tilde{\omega}_{b\rightarrow a}$ assumes a different value depending on the formulation. For instance, in BiNI [Cao2022BiNI], $\tilde{\omega}_{b\rightarrow a}$ has the following form for a pinhole camera of focal lengths $f_{x}$ and $f_{y}$ and principal point $(c_{x},c_{y})$:
@@
Milano et al[Milano2025DiscontinuityAwareNormalIntegration] instead derive the following alternative equation for generic central cameras:
@@
Figure 1: Example of surface discontinuities. Left: The chair in the foreground is separated by a full surface discontinuity from the wall in the background. Infinitely many depth maps can describe the input normal map, and solutions of the integration are up to the relative scales of the chair and the wall. Right: The surface points $\boldsymbol{p_{a}}$ and $\boldsymbol{p_{b}}$ are separated by a discontinuity along the path in red, but the yellow path connects them along the surface without discontinuities. Source of object meshes : [MeshOfficePropsLowpoly2021] (left), [Shi2016DiLiGenT] (right).
While in the general case the integration solution is up to these residuals, a number of assumptions can be made about the discontinuities. In particular, discontinuities often arise from self-occlusions rather than from fully disconnected surfaces, for instance, in normal maps that capture a single object [Shi2016DiLiGenT]. In general, even in the presence of multiple objects, while the residual $\chi_{b\rightarrow a}$ between two neighboring pixels $a$ and $b$ might be significant in magnitude, there often exists an alternative path along the surface that connects the two corresponding surface points along a trajectory with no discontinuities (Fig.˜1, right).
State-of-the-art methods implicitly exploit this fact by assigning a lower weight in the optimization objective to the constraints (1) that, over the course of the optimization, are found to have a large residual magnitude. The rationale is that if all constraints were equally weighted, those associated with discontinuities would incorrectly drive the optimization to assign them a small residual, causing the reconstruction to converge to an overly continuous surface. By iteratively adapting the weights of the constraints, these methods instead focus on the more continuous paths along the surface, and thereby more accurately recover discontinuities, while still reconstructing a connected surface. A complementary approach, recently proposed by [Milano2025DiscontinuityAwareNormalIntegration], consists in dynamically updating the coefficients $\tilde{\omega}_{b\rightarrow a}$ in the equations (1) that have large residuals, so as to explicitly take into account the magnitude of the discontinuities. This is achieved through an additive term in the logarithm (3), which admits a geometrical interpretation. We refer to the exact mechanism by which the above methods achieve the reweighting of the constraints as a model of discontinuity.
The leading model of discontinuity is the bilateral weighting scheme proposed by BiNI [Cao2022BiNI] which acts through two different mechanisms. First, it relies on the assumption that surfaces are semi-smooth. This implies that, indexing the two neighboring pixels $b$ and b-b on opposite sides of a pixel $a$, the depth map has a discontinuity between at most one of the sides $(a,b)$ or (a,b)(a,-b). This is modeled by weighting the constraint (1) between $a$ and $b$ by the factor,
@@
As noted by [Milano2025DiscontinuityAwareNormalIntegration], the factor $\gamma_{b\rightarrow a}$ can be expressed as
@@
where $f$ is the focal length of the camera, and is crucial to the optimization convergence, which it affects through (6) by introducing a multiplicative factor $(\boldsymbol{n_{a}}^{\mathsf{T}}\boldsymbol{\tau_{a}})^{2}$. Since the quantity $\boldsymbol{n_{a}}^{\mathsf{T}}\boldsymbol{\tau_{a}}$ tends to 0 for pixels $a$ that lie close to an object boundary [Marr1977AnalysisOccludingContour], this weighting effectively assigns lower confidence to constraints involving pixels that are likely to lie next to a depth discontinuity.
@@ 3.3 Formation and filling of the initial components
Importantly, since no two components share any variables or constraints, the above optimization problem can be solved independently and in parallel for each of the components. Similarly to previous log-depth-based methods [Cao2022BiNI, Milano2025DiscontinuityAwareNormalIntegration], we initialize all log-depth values to 0 and perform optimization using the conjugate-gradient method [Hestenes1952ConjugateGradient], denoted as $\mathrm{cg}(\mathbf{A},\mathbf{b})$, on the normal equations:
@@ 4.1 Experimental settings
Our implementation relies on a combination of GPU-accelerated tensor operations and CPU-based operations (e.g., conjugate-gradient-based optimization, for which we use the SciPy library [Virtanen2020SciPy]). We compare our method to the current state-of-the-art approach of [Milano2025DiscontinuityAwareNormalIntegration], noting that we also use their model of continuity in our framework. We reimplement the baseline and integrate it into our framework, to decouple the effects of GPU acceleration and formulation-specific computational aspects. All experiments are run on a single NVIDIA RTX $3080$ Laptop GPU.
@@ 4.2 Benchmark experiments
We evaluate on the DiLiGenT benchmark [Shi2016DiLiGenT] for normal integration, which features ground-truth, object-level normal maps of resolution $608\times 512$. In these maps the background is masked out, and a normal vector is defined only at the object pixels, resulting in 24 70624\,706 to 56 56056\,560 depth values to be estimated. For each object and run we report the mean average depth error (MADE) and the total execution time.
@@
For both methods we use $8$-connectivity and convergence parameters $\Delta E_{\mathrm{max}}=10^{-3}$, $T=150$. We refer the reader to the Supplementary for ablations on these parameters. Additionally, while slight speed-ups of our method can be achieved through our optional merging (see Supplementary), we find that this effect is not particularly significant, given the small number of valid pixels in the input normal maps. We therefore do not use merging in our DiLiGenT experiments, and show instead its benefit on larger-resolution maps in the next Section. We test our method for different values of the normal similarity threshold $\theta_{c}$, including the limit case in which each pixel is assigned to a component, which we denote as $\theta_{c}=\mathrm{None}$. For the baseline, we include both the version with and without discontinuity computation ($\alpha_{b\rightarrow a}$ in [Milano2025DiscontinuityAwareNormalIntegration]). Finally, we note that our proposed outlier reweighting mechanism can also be applied to the log-depth-based optimization of [Milano2025DiscontinuityAwareNormalIntegration] and we therefore evaluate both methods with and without it.
@@
The results of our evaluation are shown in Table˜1. We start by noting that our outlier reweighting significantly improves the convergence time for pixel-level approaches, not only for our relative scale optimization, but also for the log-depth-based baseline of [Milano2025DiscontinuityAwareNormalIntegration]. Importantly, for this baseline, outlier reweighting also improved accuracy, although with the exception of some objects (cf., buddha, cow, pot1). Additionally, we note that while the discontinuity computation strategy of [Milano2025DiscontinuityAwareNormalIntegration] is beneficial for objects with large discontinuities (buddha, harvest, $\mathrm{D}$ vs. $\mathrm{ND}$ in Table˜1), its effect is partially reduced by the use of outlier reweighting.
@@
Crucially, for our component-level versions, outlier reweighting is critical for achieving optimal accuracy, resulting in significantly better and more stable convergence across the board, particularly for objects with large discontinuities (buddha, harvest, reading, goblet), as can be seen by comparing successive rows in Table˜1. Overall, while with outlier reweighting the performance of the two methods is comparable for the most continuous objects, our component-based variants achieve better accuracy than log-depth optimization for the more discontinuous objects (buddha, harvest, reading). This effect is a function of the chosen value for the threshold $\theta_{c}$, with smaller values generally resulting in lower error at the cost of slightly increased runtime. We refer to the Supplementary for details on the minimum theoretical error that can be achieved for different values of $\theta_{c}$. Notably, our pixel-level version also slightly outperforms the (pixel-level) formulation of [Milano2025DiscontinuityAwareNormalIntegration], suggesting that solving for relative scales instead of log-depth may lead to a more well-posed optimization problem.
@@
Finally, we note that, when using outlier reweighting both our component-based formulation and the pixel-level method of [Milano2025DiscontinuityAwareNormalIntegration] achieve comparable runtime on the DiLiGenT benchmark. As we show in the next Section, however, the gap between the execution times of the two frameworks becomes significantly larger as the number of valid pixels in the normal map increases.
@@
[Milano2025DiscontinuityAwareNormalIntegration], $\mathrm{D}$ N/A $\mathbf{0.02}$ $2.70$ $0.62$ $3.32$ $0.05$ $2.98$ $0.13$ $1.49$ $2.07$ $6.96$ $0.47$ $\underline{3.74}$ $0.15$ $2.86$ $0.25$ $2.02$ $\mathbf{4.04}$ $3.45$
@@
[Milano2025DiscontinuityAwareNormalIntegration], $\mathrm{ND}$ N/A $\mathbf{0.02}$ $1.84$ $0.70$ $3.04$ $0.05$ $1.87$ $0.13$ $1.15$ $2.21$ $5.99$ $0.47$ $4.13$ $\underline{0.14}$ $\mathbf{1.79}$ $0.24$ $\underline{1.62}$ $\mathbf{4.04}$ $1.62$
@@
Table 1: Mean absolute depth error (MADE, abbreviated as $\mathrm{Err}$) $\mathrm{m}\mathrm{m}$] and total execution time (abbreviated as $t$) $\mathrm{s}$] on the DiLiGenT benchmark [Shi2016DiLiGenT]. For each object, bold and underlined denote respectively the best and the second-best result across the methods. All experiments use $\Delta E_{\mathrm{max}}=10^{-3}$, $T=150$, and 88-connectivity, without merging. $\mathrm{D}$ and $\mathrm{ND}$ denote the version of [Milano2025DiscontinuityAwareNormalIntegration] with and without discontinuity computation, respectively ($\alpha_{b\rightarrow a}$ in [Milano2025DiscontinuityAwareNormalIntegration]). *This object contains a full depth discontinuity.